OpenStudy (anonymous):

Calculate (1+ x+ x²+x^3+...+ x^9+x^10)²

OpenStudy (anonymous):

hint: geometric series

OpenStudy (anonymous):

another hint: 1+x+x^2 + x^3 .. + x^10 is a geometric series (1=x^0) with 11 terms what is the sum of geometric series ?

OpenStudy (anonymous):

Wait, is the formula simply x^(n-1)

OpenStudy (anonymous):

assuming x is not 1. if it is 1 so the answer is simple.

OpenStudy (anonymous):

no sum of 1+x+x^2+...+x^10 = (1-x^11) / (1-x)

OpenStudy (anonymous):

sorry for that .. it is 11 at the numerator

OpenStudy (anonymous):

so case x=1: 11^2 case x is not 1 : [ (1-x^11) / (1-x) ]^2

OpenStudy (anonymous):

squaring this is simple